# Improving Multimodal Large Language Models Using Continual Learning

### v1

@@ Abstract
Generative large language models (LLMs) exhibit impressive capabilities, which can be further augmented by integrating a pre-trained vision model into the original LLM to create a multimodal LLM (MLLM). However, this integration often significantly decreases performance on natural language understanding and generation tasks, compared to the original LLM. This study investigates this issue using the LLaVA MLLM, treating the integration as a continual learning problem. We evaluate five continual learning methods to mitigate forgetting and identify a technique that enhances visual understanding while minimizing linguistic performance loss. Our approach reduces linguistic performance degradation by up to 15% over the LLaVA recipe, while maintaining high multimodal accuracy. We also demonstrate the robustness of our method through continual learning on a sequence of vision-language tasks, effectively preserving linguistic skills while acquiring new multimodal capabilities.
Figure 1: Summary results of the best CL methods we evaluated for training LLaVA 1.5 compared to the unimodal base LLM and the original version of LLaVA 1.5. All results are with Pythia 2.8B as the base LLM. The best method has almost the same vision-language (VL) accuracy while providing a large increase in linguistic performance on 1 NLG and 4 NLU tasks by 8% and 2% (absolute), resp.
@@ 1 Introduction
Advances in integrating visual information with large language models (LLMs) have led to the development of multimodal large language models (MLLMs), excelling at many vision-language (VL) tasks [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Recent studies converge on a general recipe for developing MLLMs: Alignment of LLM token embeddings with visual embeddings followed by instruction-tuning on VL tasks like visual question answering (VQA) [8]. However, creating an MLLM often degrades the LLM’s natural language understanding (NLU) and generation (NLG) performance, a phenomenon known as catastrophic forgetting [3, 6]. For instance, PaLM-E experienced an 87% drop in NLG performance over the base LLM [6]. Similar forgetting has been noted for LLaVA [11], but little work has addressed understanding and mitigating this issue. Multimodal LLMs are designed in part to serve as general multimodal understanding models [13, 14]. They must therefore perform well not only on vision-language data, but also retain their linguistic or text-only performance.
Here, we study mitigating the loss of linguistic abilities in the popular LLaVA MLLM using continual learning (CL) techniques designed to mitigate catastrophic forgetting [15]. In CL, a sequence of non-stationary tasks is learned, where we treat the first task as already learned by the base LLM followed by new VL tasks. We study these methods in two paradigms. In the first, we seek to recreate LLaVA 1.5 while mitigating linguistic forgetting through our methods, and in the second, we sequentially learn each VL dataset in the LLaVA recipe.
@@ 2.1 The LLaVA MLLM
We study LLaVA 1.5, henceforth referred to as LLaVA, which is one of the most widely used multi-modal training protocols. LLaVA has the following components:
@@
    Visual Encoder: Following earlier implementations [16, 12], we use a pre-trained ViT-L/14 from CLIP which takes an image resolution of 336px as the vision encoder, which is kept frozen throughout training to ensure stability, prevent overfitting to initial training tasks.
@@
    LLM: We study 9 choices for LLaVA’s LLM of varying scales and instruct-tuning: 6 Pythia models (160M - 2.8B) [17], Phi2 (3B) [18], and 2 LLaMA 2 (7B) models. Refer to Appendix A.3 for details on the base LLMs.
@@
    Alignment Network: To inject other modalities into LLaVA, it uses a two layered alignment network that projects embeddings from the vision encoder into the embedding representational space of the text tokens [19].
We follow the standard LLaVA 1.5 training recipe, and provide detailed implementational descriptions in Appendix A.3.
2.2 Continual Learning Methods
To mitigate catastrophic forgetting in MLLMs, we examine and adapt several continual learning methods for LLaVA multimodal training: LoRA [20], Soft Targets [21], Rehearsal [22], and mSGM [21], along with the original LLaVA fine-tuning (Naive FT). Details of the mitigation methods, the continual learning setup and data mixtures are provided in Appendix  A.
@@ 2.2 Continual Learning Methods
In all methods, the alignment layer and LLM are trained and the ViT is frozen. Following the LLaVA 1.5 training protocol, the VL datasets are trained in a single epoch corresponding to a single training pass through each dataset.
@@ 3 Experiments
Following the standard LLaVA recipe, for all experiments, task 2 begins by training the VL alignment network using the LLaVA-595 CC-SBU captioning dataset. The network is trained to generate captions with an auto-regressive loss, with the vision encoder and LLM kept frozen. Subsequently, the LLM and alignment layer are trained for tasks 2 to $T$. For the forgetting and mitigation experiments in Secs. 3.1 and 3.2, the two-task setup is identical to LLaVA 1.5’s training protocol. The continual learning setup employed in Sec. 3.3 is detailed in Appendices  A and B.
Evaluation. We evaluate the models using six natural language datasets: Lambada [23] for NLG, ARC-Easy [24], ARC-Challenge [24], Winogrande [25], and WSC [25] for NLU, as well as four vision-language datasets: VQAv2 [26], GQA [27], TextVQA OCR and Pure [28], and RefCOCO [29] corresponding to the LLaVA data mixture (see Appendix B.3 for details). Performance is measured using accuracy and the forgetting metric $\Delta$, which quantifies changes in task performance following training on new tasks (see Appendix A.3).
@@ 3.1 Analyzing Linguistic Forgetting
Linguistic forgetting has not been studied in LLaVA, so before assessing CL methods, we measure linguistic forgetting and VL performance for all LLMs using the standard LLaVA training recipe; which uses naive fine-tuning. Using this recipe, our 9 LLMs are transformed into MLLMs.
Overall Results. Our overall results are given in Fig. 2(a). Five models suffer from linguistic forgetting, while surprisingly, four have increased NLU/NLG accuracy due to positive transfer from the VL tasks.
Positive Backwards Transfer & Analysis of NLU vs. NLG Tasks. In Table 2(b) , we study the impact of forgetting on NLU vs. NLG tasks. All MLLMs exhibit greater NLG forgetting than NLU forgetting, and NLU datasets are the source of the cases of positive backward transfer (negative $\Delta$). We posit this may be due to the additional common-sense reasoning and world knowledge encoded in visual-language tasks and instructions, which is relevant for the NLU tasks. In terms of model size, typically smaller models show higher NLG forgetting, which is consistent with PaLM-E’s observations. We include additional results on NLU, NLG forgetting in Appendix Sec. C.3.
(a) Linguistic forgetting versus VL perf. for 9 MLLMs trained with the LLaVA recipe. Five models exhibited linguistic forgetting, while four had negative linguistic forgetting, indicating that VL training resulted in positive transfer to language tasks.
@@
(b) NLU vs NLG Forgetting: Composition of linguistic forgetting between NLU and NLG tasks for models after LLaVA training. Negative $\Delta$ indicates positive backward-transfer, which is desirable. "I" denotes the instruct-tuned model.

### v2

@@ Abstract
Generative large language models (LLMs) exhibit impressive capabilities, which can be further augmented by integrating a pre-trained vision model into the original LLM to create a multimodal LLM (MLLM). However, this integration often significantly decreases performance on natural language understanding and generation tasks, compared to the original LLM. This study investigates this issue using the LLaVA MLLM, treating the integration as a continual learning problem. We evaluate five continual learning methods to mitigate forgetting and identify a technique that enhances visual understanding while minimizing linguistic performance loss. Our approach reduces linguistic performance degradation by up to 15% over the LLaVA recipe, while maintaining high multimodal accuracy. We also demonstrate the robustness of our method through continual learning on a sequence of vision-language tasks, effectively preserving linguistic skills while acquiring new multimodal capabilities. Project webpage: https://shikhar-srivastava.github.io/cl-for-improving-mllms
@@ 1 Introduction
Advances in integrating visual information with large language models (LLMs) have led to the development of multimodal large language models (MLLMs), excelling at many vision-language (VL) tasks (Wu et al., 2023; Yang et al., 2023; Li et al., 2023a; Alayrac et al., 2022; Chiang et al., 2023; Driess et al., 2023; Zhu et al., 2023a; Liu et al., 2023b; Luo et al., 2023a; Koh et al., 2023; Lin et al., 2023a; Liu et al., 2023a). Recent studies converge on a general recipe for developing MLLMs: Alignment of LLM token embeddings with visual embeddings followed by instruction-tuning on VL tasks like visual question answering (VQA) (Liu et al., 2023b). However, creating an MLLM often degrades the LLM’s natural language understanding (NLU) and generation (NLG) performance (Li et al., 2023a; Driess et al., 2023). This phenomenon of the sudden degradation in previously learned capabilities as a result of progressive training is known as catastrophic forgetting (McCloskey & Cohen, 1989). For instance, the multimodal PaLM-E experienced an 87% relative drop in performance on a range of NLG tasks compared to its unimodal base LLM (Driess et al., 2023). Recent approaches towards MLLMs, such as LLaVA, notably also suffer from the loss of linguistic abilities (Liu et al., 2023b; Lin et al., 2024). This degradation occurs primarily due to the significant distribution shift when adapting LLMs, originally trained on text-only tasks, to multimodal vision-language (VL) contexts. Limiting this degradation by freezing the LLM, training secondary modules (Dai et al., 2024; Koh et al., 2023; Zhu et al., 2023a) underperforms fully adapted approaches like LLaVA and precludes emergent multimodal capabilities like in-context learning  (Yu et al., 2023; Lin et al., 2024). Linguistic forgetting, therefore, is pervasive in the multimodal adaptation of LLMs, with limited work on studying this phenomenon and developing methods to efficiently mitigate it.
We posit that a practical MLLM must excel in both purely text-based and vision-language tasks. In real-world scenarios, users frequently alternate between text-only and multimodal queries in multi-turn interactions – interacting with multimodal systems by asking questions in text, then refining it with an image, or vice versa (Chen et al., 2024c; b). Moreover, strong linguistic understanding remains essential for performative responses even inside multimodal prompts (Lin et al., 2023a). To address these challenges, we propose employing Continual Learning (CL) approaches, explicitly developed to mitigate catastrophic forgetting arising from shifting task distributions (Parisi et al., 2019). Because the introduction of visual information constitutes a significant distributional shift, studying MLLM adaptation within a CL framework is natural, and facilitates preservation of linguistic capabilities.
Finally, a general multimodal learning system must continuously acquire new multimodal capabilities without sacrificing both pre-existing linguistic and acquired multimodal capabilities. In CL, a sequence of non-stationary tasks is learned continuously. We treat the first task as the unimodal pre-training task already learned by the base LLM, followed by new VL tasks. We study two paradigms. In the first, we seek to recreate LLaVA 1.5 while mitigating linguistic forgetting through our methods. In the second, we sequentially learn each VL dataset in the LLaVA recipe, evaluating the model’s incremental acquisition of multimodal skills without degrading previously learned capabilities.
Figure 1: Summary results of the best continual learning method we evaluated for training LLaVA 1.5 compared to the Language-Only LLM, and the original version of LLaVA 1.5. All results are with Pythia 2.8B as the base LLM. The best method has almost the same vision-language (VL) accuracy while providing a large increase in linguistic performance on 1 NLG and 4 NLU tasks by 8% and 2% (absolute), respectively. Bar annotations indicate linguistic forgetting relative to the Language-Only LLM, highlighted in red.
@@
2 Background
2.1 Multimodal Large Language Models
Open source MLLMs integrate pre-trained vision encoders into LLMs (see  Bordes et al. (2024) for review).
Notable works include PaLM-E (Driess et al., 2023), LLaVA 1.5 (Liu et al., 2023b; a), MiniGPT4 (Zhu et al., 2023a), LaVIN (Luo et al., 2023a), Visual ChatGPT (Wu et al., 2023), BLIP-2 (Li et al., 2023a), Qwen-VL (Bai et al., 2023), Flamingo (Alayrac et al., 2022), FROMAGe (Koh et al., 2023), and VILA (Lin et al., 2024). PaLM-E is an embodied MLLM that is capable of grounded reasoning for VL and language tasks. MiniGPT-4 uses a trainable fusion layer to link the LLM to a vision encoder, with both the LLM and vision encoder frozen. Similar to MiniGPT-4, Qwen-VL and Qwen-VL-Chat models use a one-layer cross-attention module to inject vision embeddings into an LLM. BLIP-2 uses a Q-Former to map visual information into the input space of LLM. LaVIN introduces Mixture-of-Modality Adaptation to extend the multimodal capabilities of LLMs with a vision encoder. FROMAGe demonstrated good performance on image captioning and image-text retrieval tasks using trainable input and output linear layers, while keeping the vision encoder and LLM frozen. In this paper, we study LLaVA, which shares many similarities with these models and is described in detail in Sec. 3.1.
While many MLLMs have been proposed, most papers focus only on VL performance and do not measure whether this resulted in any loss of NLU/NLG performance, including the original MiniGPT-4 and LLaVA papers. Notable exceptions are PaLM-E, LaVIN, and FROMAGe. However, they do not study methods for mitigating forgetting when an LLM is converted into an MLLM.  Driess et al. (2023) observed that multimodal adaptation led to catastrophic performance degradation (over 87%) on a suite of 21 general language benchmarks, a problem especially pronounced in smaller MLLMs (<<14B parameters). This is particularly concerning as smaller LLMs comprise the largest segment of open-source LLMs (Gunasekar et al., 2023; Li et al., 2023b), and are also the most widespread for consumer applications (such as on-device and edge deployments). Addressing this problem of linguistic forgetting could therefore significantly improve the utility of MLLMs. Approaches such as InstructBLIP (Dai et al., 2024), FROMAGe (Koh et al., 2023) keep the LLM and visual encoders frozen while learning secondary instruction-aware modules. These approaches limit interference in vision-language and linguistic capabilities, but have been shown to underperform LLaVA and preclude emergent multimodal capabilities like in-context learning  (Yu et al., 2023; Chen et al., 2024a; Lin et al., 2024). The only work that has attempted to address NLU/NLG forgetting in a comparatively performative MLLM is VILA (Lin et al., 2024), which did so by including MMC4 in their training recipe, which is a dataset that interleaves text and images (Zhu et al., 2023b). Training for this interleaved image-text task also required a large increase in pre-training compute. Moreover, they only study its impact on a single NLU dataset, neglecting its potential impact on NLG tasks. Our work is the first to study mitigating linguistic forgetting in MLLMs using CL methods.
2.2 Continual Learning
Creating an MLLM maps directly onto a CL scenario involving learning two tasks sequentially, where the first task is acquiring NLU/NLG capabilities and the second is acquiring VL capabilities. Although some recent works have studied CL with unimodal LLMs (Luo et al., 2023b; Lin et al., 2023b), our work is the first to treat creating an MLLM as a CL problem. While there are many CL techniques, two general families of methods are most appropriate for MLLMs: rehearsal and regularization. Rehearsal methods store or reconstruct previously observed data that is mixed into new data to overcome catastrophic forgetting (Hayes et al., 2021; 2020; Chaudhry et al., 2019; Harun et al., 2023; Hou et al., 2019; Rebuffi et al., 2017). The primary challenge with using it for MLLMs is that the training sets for many open-source LLMs have not been released, and even if they were released, determining what to rehearse may be intractable due to the large size of these datasets. Regularization methods adjust training processes and impose constraints on weight updates to reduce forgetting (Chaudhry et al., 2018b; a; Dhar et al., 2019; Kirkpatrick et al., 2017; Zenke et al., 2017). We focus on lightweight regularization methods in this work that can scale to MLLMs.
Parameter Efficient Tuning (PET) methods are an efficient alternative to full fine-tuning (Houlsby et al., 2019). They update fewer parameters, thereby reducing computational costs while maintaining performance. In Harun & Kanan (2024), LoRA was shown to effectively retain performance in ImageNet pre-trained models as new classes were continually learned. Inspired by this, we study the effectiveness of LoRA (Hu et al., 2021) in mitigating catastrophic forgetting while learning VL tasks.
@@ 3.1 The LLaVA MLLM
We study LLaVA 1.5, henceforth referred to as LLaVA, which is one of the most widely used multimodal training protocols. LLaVA has the following components:
  • Visual Encoder: Following earlier implementations (Liu et al., 2024; 2023a), we use a pre-trained ViT-L/14 from CLIP which takes an image resolution of 336px as the vision encoder, which is kept frozen throughout the training to ensure stability and prevent overfitting to the initial training tasks.
  • LLM: We study 9 LLMs of various sizes and training protocols (see Sec. 3.1.1).
  • Alignment Network: To inject other modalities into LLaVA, it uses a linear alignment network that projects embeddings from vision into the representational space of the text tokens (Merullo et al., 2022). This network has two layers that map each vision embedding to the text embedding space.
We follow the standard LLaVA 1.5 training recipe, with details, including optimizer settings and hardware, provided in Section 3.4 and Appendix  B.
3.1.1 LLMs Studied
We study 9 choices for LLaVA’s LLM: 6 Pythia models (Biderman et al., 2023), Phi2 (3B) (Gunasekar et al., 2023), and 2 LLaMA 2 (7B) models. The 6 Pythia models are of various scales (160M, 410M, 1B, 1.4B, and 2.8B parameters), where there are 2 versions of Pythia (1.4B), one of which is instruction fine-tuned. Likewise, we study the original version of LLaMA 2 and an instruction fine-tuned version, Vicuna-1.5 7B (Chiang et al., 2023), which was used in LLaVA 1.5.
Phi2 (Li et al., 2023b) has been trained on the same dataset as Phi1 (Gunasekar et al., 2023), which includes a curated selection of “textbook quality” data from the web (6 billion tokens) and an additional 1 billion tokens of synthetically generated textbooks and exercises created using GPT-3.5 (Brown et al., 2020). The Phi series contain performant models given their small parameter size class such as Phi2 with 2.7B parameters (Li et al., 2023b). Pythia (Biderman et al., 2023) comprises two sets of 8 models, each corresponding to two datasets. For every model size, one set is trained on the Pile dataset (Gao et al., 2020), while the other set is trained on a version of the Pile where global de-duplication has been applied. The granular model scaling suite of Pythia is particularly useful for our study. We select the de-duplicated set of Pythia models. LLaMA 2 is reportedly trained on a mix of publicly available online data, but specific details are not available (Touvron et al., 2023). LLaVA 1.5 reports its best performance with instruction fine-tuned LLaMA 2 LLMs (Liu et al., 2023a).
3.2 Continual Learning Methods
To mitigate catastrophic forgetting in MLLMs, we examine several methods:
@@ 3.2 Continual Learning Methods
    LoRA keeps the original LLM weights frozen and learns low-rank updates for them (Hu et al., 2021). Following LoRA’s recommended protocol, we inject LoRA weights into all LLM linear layers. After each task is learned, the LoRA weights are merged into the LLM. Details are given in Appendix B.1.2.
@@
    Soft Targets, introduced in Harun & Kanan (2024), aim to mitigate forgetting in CL. However, they are designed for image classification tasks and are not directly applicable to MLLMs. We adapt soft targets to MLLMs by dynamically re-parameterizing the conditional token distribution. We update the hard target vector $Y$ by reducing the target tokens by $\alpha$, and offsetting non-target tokens by $\alpha/(N-1)$, where $\alpha$ controls the smoothing, and $N$ is the LLM’s vocabulary size. We discuss the choice of $\alpha$ in Appendix B.1.3.
  4. 4.
    Rehearsal (Experience Replay) is an effective method for CL that involves storing data from earlier tasks and mixing it with data from current task. We study it in our CL experiments since in practice, the pretraining data for the base LLMs (i.e. task 1) are inaccessible. We study storing 1% of randomly selected samples from each previous task, excluding task 1.
  5. 5.
    mSGM is based on SGM, which combines soft targets, weight initialization, and LoRA to mitigate catastrophic forgetting (Harun & Kanan, 2024). We use a version that omits the weight initialization and output layer freezing process used by SGM for class incremental learning since the output vocabulary is static and output layers are used for the causal generation of LLMs. Harun & Kanan (2024) found that SGM was especially effective when paired with rehearsal.
@@
In all methods, the alignment layer and LLM are trained and the ViT is frozen. Following the LLaVA 1.5 training protocol, the VL datasets are trained in a single epoch corresponding to a single training pass through each dataset after learning task 1 to update the LLM.
3.3 Evaluation Datasets
To facilitate analysis across tasks, all of our evaluation sets use the same metric: accuracy.
Natural Language Evaluation: We use six datasets. For NLG, we use Lambada (Paperno et al., 2016), which is the only NLG dataset that uses accuracy for evaluation. For NLU, we use ARC-Easy (Clark et al., 2018), ARC-Challenge (Clark et al., 2018), Winogrande (Sakaguchi et al., 2021), and WSC (Sakaguchi et al., 2021) for NLU. For completeness, we also study multiple NLG datasets using a ‘% Change’ $\Delta^{\prime}$ metric in Appendix D.2.
Vision-Language Evaluation: We use the test sets corresponding to each VL dataset used for training LLaVA: VQAv2 and GQA for general VQA tasks (Goyal et al., 2017; Hudson & Manning, 2019), TextVQA OCR (and Pure) for OCR tasks (Singh et al., 2019), and RefCOCO for referential expression generation tasks (Kazemzadeh et al., 2014). We use the slim version of all test datasets to make our experiments tractable  [^1] https://github.com/TRI-ML/vlm-evaluation. Additional details are given in Appendix C.
3.4 LLaVA Training Details
Following the LLaVA 1.5 protocol  (Liu et al., 2023b), the visual encoder is a CLIP VIT-L@336px with an image resolution of 336px with letterbox resizing, and the alignment network is a two-hidden layer MLP projector with GELU activation. The LLM generations are limited to a maximum token length of $2048$. For training, we used the Adam optimizer, with FSDP (Fully Sharded Data Parallel), gradient checkpointing and mixed precision to train all our models (Zhao et al., 2023).
Alignment Stage: During LLaVA 1.5’s alignment stage the LLM is kept frozen, while the alignment network is plastic. We use a learning rate of $0.001$ with a linear warmup followed by a cosine decay scheduler. The training process is carried out for $1$ epoch with a global batch size of $256$, distributed as $16$ samples per device. During the alignment stage, we shard only the gradients and optimizer states, not the parameters.
Fine-tuning Stage: During LLaVA 1.5’s fine-tuning stage both the LLM and alignment networks are jointly fine-tuned. We set the learning rate to 2e052e-05, with a linear warmup followed by cosine decay. This stage also runs for 1 epoch but with a smaller global batch size of $128$, keeping the per-device batch size constant at $16$. Depending on the size of the LLM (Phi2 3B vs Pythia 160M), we vary the per-device batch size, but the global batch size is kept constant across all experiments. We switch to the FSDP full sharding strategy, with all parameters, optimizer states, and gradients sharded across the devices.
Continual LLaVA: For the Continual LLaVA setting, we follow the same experimental configuration as before in the LLaVA setting. A model trained on Task $(i)$, will then be simply used as the pretrained checkpoint for Task $(i+1)$, with training following normally. The training configurations are then identical to the fine-tuning stage. In the case of PET methods like LoRA, the adapter weights are merged back into the LLM before starting training on the new task.
To reproduce the LLaVA 1.5 experiment, we build on top of the Prismatic library (Karamcheti et al., 2024), and use the Prismatic library for vision-language tasks and EleutherAI’s LM_Eval library (Gao et al., 2023) for natural language evaluations. All our code is written in PyTorch (Paszke et al., 2019).
3.5 Measuring Performance
To assess performance, we compute the forgetting metric $\Delta$, where a positive value indicates forgetting, and a negative value indicates learning the current task enhances the performance of previously acquired tasks (or a backward transfer (Zenke et al., 2017)). The performance change, $\Delta$, on task $t$ after training on task $k$ is defined as:
$\Delta_{t}(k)=\omega_{t}(1)-\omega_{t}(k),\quad\forall t<T$ (1)
where $T$ is the total number of tasks, and $\omega_{t}(k)$ is the harmonic mean of the accuracy values of the evaluation datasets for task $t$, after training on task $k$. We use the harmonic mean because it is dominated by the accuracy of the worst-performing dataset for the task, thereby emphasizing the need to perform well, avoid forgetting for all datasets.
@@ 4 Experiments
Following the standard LLaVA recipe, for all experiments, task 2 begins by training the VL alignment network using the LLaVA-595 CC-SBU captioning dataset. The network is trained to generate captions with an auto-regressive loss, with the vision encoder and LLM kept frozen. Because only the alignment network is trained, the model has a modicum of visual ability without impairing its language capabilities. Subsequently, the LLM and alignment layer are trained for tasks 2 to $T$. For Secs. 4.1 and 4.2, their two-task setup is identical to LLaVA 1.5’s training protocol. Additional implementation details are given in Sec. 3.4.
@@ 4.1 Analyzing Linguistic Forgetting
Linguistic forgetting has not been studied in LLaVA, so before assessing CL methods, we first measure linguistic forgetting and VL performance for all of our LLMs using the standard LLaVA training recipe; which uses naive fine-tuning. Using this recipe, our 9 LLMs are transformed into MLLMs.
Figure 2: Linguistic forgetting vs. VL performance for 9 MLLMs trained with the LLaVA recipe. Five models exhibited linguistic forgetting, and four had negative linguistic forgetting values, indicating that VL training resulted in positive transfer to language tasks.
Overall Results.
Our overall results are given in Fig. 2. Five models suffer from linguistic forgetting, while surprisingly, four have increased NLU/NLG accuracy due to positive transfer from the VL tasks. We next dive into these results, where our analysis of NLU/NLG forgetting as a function of model scale is provided in a later subsection.
Positive Backwards Transfer & Analysis of NLU vs. NLG Tasks.
In Table 2, we study the impact of forgetting on NLU vs. NLG tasks. We find that forgetting is much more severe for our one NLG task than the NLU tasks. All MLLMs exhibit greater NLG forgetting than NLU forgetting as shown in Table 2. In terms of model size, typically smaller models show higher NLG forgetting, which is consistent with PaLM-E’s observations. Across MLLMs, NLG forgetting is much more severe compared to NLU, and the NLU datasets are the source of the cases of positive backward transfer (negative $\Delta$). We posit this may be due to the additional common-sense reasoning and world knowledge encoded in visual-language tasks and instructions, which is relevant for the NLU tasks. We include additional results on the NLU and NLG forgetting in Appendix D.1 (see Fig. 5), and analyze linguistic generation capabilities across multiple NLG benchmarks in Appendix D.2.
@@
Table 1: NLU vs NLG Forgetting: Negative $\Delta$ indicates positive backward-transfer (desirable). The letter (I) denotes an instruction-tuned model.
Base LLM Instr. VL Avg. NL Avg.
Acc $\uparrow$ Acc $\uparrow$ $\Delta\downarrow$
LLAMA 2 (7B) 57.22 66.23  –1.84
56.55 64.44  –0.36
Pythia (1.4B) 43.97 45.51    2.18
43.93 41.37  –1.16
Table 2: Results for LLaMA 2 and Pythia after LLaVA training, with and without instruction fine-tuning. Negative $\Delta$ indicates positive backward transfer.
The Impact of Instruction Fine-Tuning.
Previous work on continual learning in large language models (LLMs) has shown that instruction fine-tuning mitigates catastrophic forgetting (Luo et al., 2023b). This study extends the investigation to multimodal language models (MLLMs), analyzing LLaVA models trained with Pythia (1.4B) and LLaMA 2 (7B), and their instruction fine-tuned versions: Vicuna 1.5, used in the original LLaVA paper (Liu et al., 2023a), and an instruction-tuned Pythia (1.4B) [^2] https://huggingface.co/lambdalabs/pythia-1.4b-deduped-synthetic-instruct. As shown in Table 2, non-instruction-tuned LLMs exhibit minor increases in visual-linguistic (VL) accuracy (<1<1%) and a larger difference in natural language understanding/generation (NLU/NLG) accuracy compared to their instruction-tuned counterparts. The individual task accuracy breakups are shown in Tables 11,  12 in the Appendix. Instruction-tuned LLaMA 2 (7B) and Pythia (1.4B) models show minimal or even negative linguistic forgetting, indicating positive backward transfer. Conversely, base LLMs display greater variability, with base Pythia experiencing more forgetting than base LLaMA 2, possibly due to differences in their pre-training datasets affecting post-multimodal training linguistic performance. Instruction tuning significantly reduces NLG forgetting in Pythia (1.4B) and induces positive backward transfer from VL to NLG, with a similar trend observed in NLU tasks. This suggests that instruction-tuned LLMs exhibit little to no forgetting, likely due to the distributional similarity between text-only instruction tuning tasks and visual-instruction tuning, helping to mitigate catastrophic forgetting.